A current of 4.91 A is passed through a Fe(NO3)2 solution. How long (in hours) would this current have to be applied to plate out 7.60 g of iron?
Answer #1

Atomic mass of iron = 56gm/mol

Equivalent weight of Fe = atomic mass /valency = 56 /2 = 28 gm/equiv

According to Faradays 1st law of electrolysis,

mass of substance deposited, W = Zct

Z= electro echemical equivalent= Equivalent weight / 96500 = 28 / 96500 = 0.0003

c= current in amp = 4.91 A

t= time in sec

W= 7.60gm

Thus, t = W /Zc = 7.60 /(0.0003 x 4.91 ) = 5159.538 5160 sec = 1.43 hr

Thus,  current have to be applied for 1.43 hour to plate out 7.60 g of iron